OpenStudy (anonymous):

Classify the graph of the equation as a circle, a parabola, an ellipse, or a hyperbola. 2x^2 + 6y^2 + 5x = 0 A. circle B. parabola C. ellipse D. hyperbola

OpenStudy (anonymous):

ellipse.

OpenStudy (anonymous):

general eq of conic is \[ax^2+ by^2+2gx+afy+ahxy+c=0\]

OpenStudy (anonymous):

conic depends on the sign of \[h^2- ab\]

OpenStudy (anonymous):

ok so negative makes it an ellipse.? 6x2 – 3y2 + 5x – 2y – 7 = 0 would be a hyperbola right or no cause it's negative

OpenStudy (anonymous):

2fy*

OpenStudy (anonymous):

yes its an ellipse

OpenStudy (anonymous):

+ve makes a hyperbola

OpenStudy (anonymous):

ok so whats makes something a parabola?

OpenStudy (anonymous):

=zero

OpenStudy (amistre64):

a parabola bends once; onlyone variable will be squared

OpenStudy (anonymous):

oh ok thanks

OpenStudy (anonymous):

I guess I just dont get how you tell the difference between a hyperbola and ellipse if they are both negative, like 5x2 – 8y2 + 7x + 3y – 7 = 0 it's not a parabola or circle

OpenStudy (amistre64):

that looks like a hyperbola

OpenStudy (anonymous):

how do you tell the difference between a hyperbola and ellipse

OpenStudy (amistre64):

when x^2 or y^2; only one of them is negative; its a hyperbola usually

OpenStudy (anonymous):

if h^2-ab is positive, the conic is a hyperbola

OpenStudy (amistre64):

an ellipse/cirlce has the same sign on x^2 and y^2

OpenStudy (anonymous):

yes if a n b possess different sign it is usualy a hyperbola

OpenStudy (amistre64):

a circle is just a special ellipse

OpenStudy (anonymous):

oh ok I get it now thanks very much

OpenStudy (amistre64):

just like a square is a special rectangle

OpenStudy (anonymous):

right

OpenStudy (amistre64):

a hyperbola is an inside out ellipse

OpenStudy (anonymous):

oh ok good way to put it